OpenStudy (anonymous):

A wooden pyramid, 6 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 3 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids.

OpenStudy (campbell_st):

the ratio of the linear measurements is 1:3 the ratio of volumes can be found by raising each value in the ratio to the power of 3 (1)^3 :(3)^3 = 1:27 a test.... height 6 and assume the square base has side length of 1 inch \[V = 1/3 \times 1^2 \times 6\] V = 2 inches^3 now with the dimensions tripled h = 18 and square side length is 3 \[V = 1/3 \times 3^2\times 18\] V = 54 inches ^3 the ratio of the volumes is 2:54 or 1: 27

OpenStudy (anonymous):

Thanks!